OpenStudy (anonymous):

Help! Circle C has a radius of 5cm Find the length of arc AB Pic below

OpenStudy (anonymous):



OpenStudy (ash2326):

Length of Arc if the angle is 360 degree or 2pi is the circumference \[Length\ of\ arc=2\pi\times r\] So arc length =angles * r If the angle is 36, what would be the arc length? Can you try now?

OpenStudy (anonymous):

length of arc=31.42

OpenStudy (anonymous):

\[A _{L}=r \times \theta \] Al=Arc length r=radius theta = angle in radian.....

OpenStudy (ash2326):

@heather040200 yes you're right, good work :)

OpenStudy (anonymous):

thank you very much for your help

OpenStudy (anonymous):

you can use this formula to also find the arc length http://math.about.com/od/formulas/ss/surfaceareavol_9.htm